OpenStudy (anonymous):

Which of the following is the slope between the two points (1, 5) and (6, 4)? 1 over 5 negative 1 over 5 5 -5 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

use the slope formula m = (y2 - y1)/(x2 - x1)

jimthompson5910 (jim_thompson5910):

In this case x1 = 1 y1 = 5 x2 = 6 y2 = 4

OpenStudy (anonymous):

answer is -1/5

OpenStudy (anonymous):

http://sketchtoy.com/32769295 Like this @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

x2 doesn't mean x^2 x2 is just the second x coordinate

OpenStudy (anonymous):

I know, I just did that to mark it.

jimthompson5910 (jim_thompson5910):

m = (y2 - y1)/(x2 - x1) m = (4-5)/(6-1) m = -1/5

jimthompson5910 (jim_thompson5910):

oh then you would mark the 2 below the x like this 

OpenStudy (anonymous):

Thanks again!

jimthompson5910 (jim_thompson5910):

np